Calculate the pH of a solution obtained by mixing 0.1 litre of pH = 4 and 0.2 litre of pH = 10

Grade:10

1 Answers

Arun
6 years ago
 

ph of strong acid = 4

 

[H+] = 10-4 = molarity of acid

 

in 0.1L , moles of acid = 0.1* 10-4

 

ph of base =10

 

[H+]=10-10

 

[OH-]=10-14/10-10=10-4 = molarity of base

 

in 0.2L , moles of acid = 0.2*10-4

 

now nutralization will occur and acid is completely nutralised by base , at the end base remaining due to higher concentration....

 

now final molariry = remaining moles/total vol = (0.2*10-4-0.1*10-4 ) /0.3

 

                     concentration of base               =10-4/3

 

[H+] = 3*10-14/10-4 =3*10-10

 

ph=-loh[H+] = 10-log3= 9.523